# Parallel Digital Twin-driven Deep Reinforcement Learning for User Association and Load Balancing in Dynamic Wireless Networks

### v1

@@ V-A1 Forward Diffusion Process
$q\left(\boldsymbol{l}_{1:T}|\boldsymbol{l}_{0}\right)=\prod_{t=1}^{T}q\left(\boldsymbol{l}_{t}|\boldsymbol{l}_{t-1}\right),$ (13)
where $\alpha_{t}\text{ for }t=1,2,\dots,T$ are hyper-parameters of the noising process, $\mathcal{N}(x;\mu,\sigma)$ represents the normal distribution of mean $\mu$ and covariance $\sigma$ that produces $x$, and $\mathbf{I}$ denotes the identity matrix. This forward process noises a real trajectory $\boldsymbol{l}_{0}$ into $\boldsymbol{l}_{T}$, making it indistinguishable from Gaussian noise at the $T$-th step. Using the properties of Gaussian distributions, the forward process can also be marginalized to directly obtain the noised trajectory at the $t$-th step. That is,
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. Additionally, the parameterization of the Gaussian distribution of the forward process allows a closed-form formulation of the posterior distribution of $\boldsymbol{l}_{t-1}$ conditioned on $\boldsymbol{l}_{0}$ and $\boldsymbol{l}_{t}$. It follows
@@ V-A2 Reverse Diffusion Process
$p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{0:T}\right)=p\left(\boldsymbol{l}_{T}\right)\prod_{t=1}^{T}p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{t-1}|\boldsymbol{l}_{t}\right).$ (18)
@@
$p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{0:T}|\boldsymbol{m}\right)=p\left(\boldsymbol{l}_{T}\right)\prod_{t=1}^{T}p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{t-1}|\boldsymbol{l}_{t},\boldsymbol{m}\right).$ (20)
@@ V-A3 Training of Map2Traj
$\min_{\boldsymbol{\theta}}\ \mathbb{E}_{\boldsymbol{m},\boldsymbol{l},t,\boldsymbol{\epsilon}}\|f_{\boldsymbol{\theta}}(\boldsymbol{m},\underbrace{\sqrt{\gamma_{t}}\boldsymbol{l}_{0}+\sqrt{1-\gamma_{t}}\boldsymbol{\epsilon}}_{\boldsymbol{l}_{t}},t)-\boldsymbol{\epsilon}\|^{2}.$ (22)
@@ V-A4 Inference of Map2Traj
The sampling process of the diffusion model starts with pure Gaussian noise $\boldsymbol{l}_{T}$, followed by $T$ refinement steps. Given any noisy trajectory $\boldsymbol{l}_{t}$, we can approximate the target trajectory by rearranging the terms in (21) as
$\hat{\boldsymbol{l}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\boldsymbol{l}_{t}-\sqrt{1-\gamma_{t}}f_{\boldsymbol{\theta}}\left(\boldsymbol{m},\boldsymbol{l}_{t},t\right)\right).$ (23)
@@
V-A5 Neural Network Model in Map2Traj
The neural network model utilized in Map2Traj, $f_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{l}_{t},t)$, is based on a 192$\times$192 U-Net model [38], with multiple modifications to improve its performance such as attention blocks [39] and group normalization [40]. A distinctive feature of Map2Traj is the incorporation of street map data through concatenation, allowing the model to be conditioned on the spatial information inherent in the maps. As illustrated in Fig. 5, street map $\boldsymbol{m}$ is concatenated with the noisy trajectory image $\boldsymbol{l}_{t}$, composing a multi-channel image as the input of the U-Net. Through multiple convolution, attention mechanism, downsampling, and upsampling blocks, this multi-channel image is transformed into the prediction of the noise component in $\boldsymbol{l}_{t}$. The prediction, i.e., $f_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{l}_{t},t)$, is represented as a single-channel image. The embedded time step $t$ is introduced to help the model assess the noise level in the trajectory image. We also provide an exemplified demonstration of the denoising process in (25) at the lower part of Fig. 5.
@@ V-A5 Neural Network Model in Map2Traj
Now we can construct the DT for the wireless network in the area of interest, with the wireless channel model based on measurement and ray-tracing, as well as the user mobility model based on Map2Traj. We first place BSs in the DT with the same number, locations, power, and bandwidth as in the real network. This ensures that the state and action spaces remain statistically identical. Then the wireless channel model is used to compute the RSRP between the BSs and users, and the user mobility model generates unique trajectories for each user in the environment. As users keep entering and leaving this area, the total number of users, $|\mathcal{U}|$, evolves over a long-time scale.
@@ V-B Parallel DT Framework
Specifically, the evolution of $|\mathcal{U}|$ in an environment occurs over a long-time scale (hundreds of time slots), while sample collection and batch training of $\pi_{\boldsymbol{\theta}}$ are performed over just one or a few time slots. As a result, the number of users in each training batch remains relatively stable or fluctuates slightly. This strong correlation of training samples on $|\mathcal{U}|$ would cause non-stationarity in DRL training, resulting in a suboptimal policy that focuses primarily on the current state, as schematized in Fig. 6
@@
One potential solution to this issue in a single environment is to extend both training interval and batch size, ensuring the batch training occurs only after $|\mathcal{U}|$ has sufficiently evolved. However, this approach would significantly reduce training efficiency. In multi-agent DRL systems, $\pi_{\boldsymbol{\theta}}$ can only undergo a limited number of gradient descents before being redeployed for further exploration. This is because the state and transition probabilities of MDP for each agent are significantly affected by actions of other agents. In the environment studied in this paper, this effect is observed in the strong relationship between actions of other agents under policy $\pi_{\boldsymbol{\theta}}$ and the distribution and transition probabilities of $\boldsymbol{l}[n]$ within the state $\mathbf{s}_{i}[n]$. While extending the training interval can mitigate this correlation, only a small portion of the collected samples would be effectively utilized for training, thus reducing the training efficiency. For instance, in our experiments, the traversal of the number of users from 100 to 400 requires 20,000 environment steps, generating $5\times 10^{6}$ training samples which are sufficient for DRL agents to converge. However, with an extended training interval, these samples would only support a few gradient descents, resulting in most of the samples being wasted. This low training efficiency is impractical for DRL training in both real and DT environments.
@@
In summary, the proposed parallel DT-driven DRL method consists of the distributed DRL method, the high-fidelity DT construction technique, and the parallel DT framework. It addresses high trial-and-error costs and poor physical network performance by training in virtual DT environments. It also alleviates the strong correlation and non-stationarity in the real environment through parallelization, which therefore improves training efficiency and the generalization ability of DRL models. Note that the DT construction technique and parallel DT framework are versatile and applicable for other DT-driven network optimization tasks, particularly mobility-sensitive ones, such as wireless resource management [41] and task offloading in mobile edge computing [42].
@@ VI Performance Evaluation
In this section, we assess the efficacy of our proposed parallel DT-driven DRL method. First, we examine the established DT’s fidelity to the real environment and its reliability in the DRL training. Then a comprehensive evaluation of the proposed parallel DT-DRL method is conducted, along with a comparison against the DRL method trained directly in the real environment.
@@ VI-B Dataset and Experiment setting
Map2Traj is trained using a real-world vehicle trajectory dataset from Xi’an, China, recorded in 2016 [43], alongside the OpenStreetMap dataset of the same year. These datasets are sourced from the ChinaGEOSS Data Sharing Network [^2] https://chinageoss.cn/. The trajectories are all within the latitude range of 34.21 to 34.28 and the longitude range of 108.912 to 108.996. To avoid data leakage in the zero-shot trajectory generation, the training set is limited to longitudes between 108.912 and 108.974, while the test set extends from 108.974 to 108.996. Notice that the correlation between street map and trajectory, as shown in Fig. 3, remains consistent under transformations such as rotation and reflection. This inherent property can be leveraged for data augmentation. We randomly rotate and flip both street maps and trajectory data in training to enhance the generalization capability of the Map2Traj model.
@@ VI-C Evaluation Metrics
The metrics for spatial distribution consist of cosine similarity and Wasserstein distance [48]. Cosine similarity is a widely used measure of similarity between two vectors. While it reflects the similarity between probability distributions, it falls short in expressing the spatial correlation between adjacent blocks in two-dimensional (2D) distributions. To address this limitation, we introduce the Wasserstein distance, which quantifies the minimal cost required to transform one probability distribution into another through an optimal transport plan. In this experiment, it measures the effort required to transform the spatial distribution of generated trajectories into that of real trajectories. However, computing Wasserstein distance for 2D distributions entails solving a complex high-dimensional linear programming problem, which is computationally impractical for 192$\times$192 arrays. Consequently, we use sliced Wasserstein distance as an alternative [49].
@@ VI-D Numerical Results
We select the area mentioned in V-A as the test area. Fig. 9 displays the generated trajectories by all methods alongside corresponding heatmaps. Traditional random mobility models, i.e., RWP and GM, result in chaotic trajectories and heatmaps that bear no resemblance to real-world patterns. While map-restricted models, M-RWP and M-GM, show some similarity in trajectory patterns to real ones, they still fall short in spatial distribution similarity due to the absence of a learning mechanism. As expected, DiffTraj demonstrates high similarity to real trajectories and heatmaps because it is trained directly on real data. In contrast, our proposed Map2Traj, even in a zero-shot scenario, produces results comparable to real trajectories and even surpasses DiffTraj in trajectory details.
(a) Real street map
(b) Real trajectories
(c) Real heatmap
(d) RWP
(e) GM
(f) M-RWP
(g) M-GM
(h) DiffTraj
(i) Map2Traj
@@
It is encouraging to see our zero-shot Map2Traj model outperforms the area-specific DiffTraj. The advantage of Map2Traj lies in the continuous guidance from street maps throughout the denoising process, while DiffTraj only involves some trip information as the condition. On the other hand, DiffTraj fixes the trajectories into uniform shapes through sampling, instead of transforming them into images, potentially leading to information loss. However, it is crucial to acknowledge that existing learning-based methods, including DiffTraj, have the potential to outmatch Map2Traj given a sufficiently large training dataset, a wider and deeper network structure, or in some specific test scenarios. The primary contribution of Map2Traj is the zero-shot trajectory generation ability, rather than merely surpassing existing area-specific trajectory generation techniques.
Figure 10: DRL performance in the Map2Traj-based training environment.
@@
Figure 11: DRL performance in real environment.
@@
As illustrated in Fig. 12, both similarities in spatial distribution and trajectory pattern are positively correlated to the performance of DRL agents. This finding verifies our analysis of the prerequisite for constructing faithful DT for network optimization, as discussed in Section V-A.
@@
Figure 14: DRL performance in training environment
@@
Figure 15: DRL performance in real environment
@@ VII Conclusion
In this paper, we have proposed a parallel DT-driven DRL method designed for user association and load balancing in dynamic wireless networks. To accommodate varying user numbers, we have introduced a distributed DRL strategy with an enhanced PPO network structure to speed up convergence. To address high trial-and-error costs in real-world training and poor physical network performance before convergence, we have developed a high-fidelity DT construction method, featuring a zero-shot generative user mobility model named Map2Traj. DRL agents undergo training within this DT environment, thereby avoiding direct interaction with the physical network. Furthermore, a parallel DT framework has been developed to alleviate the strong correlation and non-stationarity in the real environment through parallelization, which therefore improves training efficiency and the generalization ability of DRL models. The proposed parallel DT-driven DRL method is versatile and applicable for other DT-driven network optimization tasks, such as wireless resource management and task offloading in mobile edge computing.

### v2

@@ V-A1 Forward Diffusion Process
$q\left(\boldsymbol{l}_{1:T}|\boldsymbol{l}_{0}\right)=\prod\nolimits_{t=1}^{T}q\left(\boldsymbol{l}_{t}|\boldsymbol{l}_{t-1}\right),$ (12)
where $\alpha_{t}\text{ for }t=1,2,\dots,T$ are hyper-parameters of the noising process, $\mathcal{N}(x;\mu,\sigma)$ represents the normal distribution of mean $\mu$ and covariance $\sigma$ that produces $x$, and $\mathbf{I}$ denotes the identity matrix. This forward process noises a real trajectory $\boldsymbol{l}_{0}$ into $\boldsymbol{l}_{T}$, making it indistinguishable from Gaussian noise at the $T$-th step. Using the properties of Gaussian distributions, the forward process can also be marginalized to directly obtain the noisy trajectory at the $t$-th step. That is,
@@
where $\gamma_{t}=\prod\nolimits_{i=1}^{t}\alpha_{i}$. Additionally, the parameterization of the Gaussian distribution of the forward process allows a closed-form formulation of the posterior distribution of $\boldsymbol{l}_{t-1}$ conditioned on $\boldsymbol{l}_{0}$ and $\boldsymbol{l}_{t}$. It follows
@@ V-A2 Reverse Diffusion Process
$p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{0:T}\right)=p\left(\boldsymbol{l}_{T}\right)\prod\nolimits_{t=1}^{T}p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{t-1}|\boldsymbol{l}_{t}\right).$ (17)
@@
$p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{0:T}|\boldsymbol{m}\right)=p\left(\boldsymbol{l}_{T}\right)\prod\nolimits_{t=1}^{T}p_{\boldsymbol{\theta}}\left(\boldsymbol{l}_{t-1}|\boldsymbol{l}_{t},\boldsymbol{m}\right).$ (19)
@@ V-A3 Training of Map2Traj
$\min\nolimits_{\boldsymbol{\theta}}\ \mathbb{E}_{\boldsymbol{m},\boldsymbol{l},t,\boldsymbol{\epsilon}}\|f_{\boldsymbol{\theta}}(\boldsymbol{m},\sqrt{\gamma_{t}}\boldsymbol{l}_{0}+\sqrt{1-\gamma_{t}}\boldsymbol{\epsilon},t)-\boldsymbol{\epsilon}\|^{2}.$ (21)
@@ V-A4 Inference of Map2Traj
The sampling process starts with pure Gaussian noise $\boldsymbol{l}_{T}$, followed by $T$ refinement steps. Given any noisy trajectory $\boldsymbol{l}_{t}$, we can approximate the target trajectory by rearranging terms in (20) as
$\hat{\boldsymbol{l}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\big{(}\boldsymbol{l}_{t}-\sqrt{1-\gamma_{t}}f_{\boldsymbol{\theta}}\left(\boldsymbol{m},\boldsymbol{l}_{t},t\right)\big{)}.$ (22)
@@
V-A5 Post Processing of Trajectory Image
To model the user mobility in wireless network optimization, the trajectory sequence needs to be reconstructed from the image. Since each image represents only one trajectory per user, we used a simple search algorithm to reconstruct the sequence from one end to the other, where each point represents the coordinate of the pixel center. An extra deep neural network is utilized to estimate the sojourn time at each point, as employed in [42], from which the speed and relative time stamp are derived.
@@ V-A5 Post Processing of Trajectory Image
V-A6 Neural Network Model in Map2Traj
The neural network model utilized in Map2Traj, $f_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{l}_{t},t)$, is based on a 192$\times$192 U-Net model [43], with multiple modifications to improve its performance such as attention blocks [44] and group normalization [45]. A distinctive feature of Map2Traj is the incorporation of street map data through concatenation, allowing the model to be conditioned on the spatial information inherent in the maps. As illustrated in Fig. 4, street map $\boldsymbol{m}$ is concatenated with the noisy trajectory image $\boldsymbol{l}_{t}$, composing a multi-channel image as the input of the U-Net. Through multiple convolutions, attention mechanism, downsampling, and upsampling blocks, this multi-channel image is transformed into the prediction of the noise component in $\boldsymbol{l}_{t}$. The prediction, i.e., $f_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{l}_{t},t)$, is represented as a single-channel image. The embedded time step $t$ is introduced to help the model assess the noise level in the trajectory image. We also provide an exemplified demonstration of the denoising process in (24) at the lower part of Fig. 4.
Now we can construct the DT for the wireless network in the area of interest, with the wireless channel model based on measurement and ray-tracing, as well as the user mobility model based on Map2Traj. We first place BSs in the DT with the same number, locations, power, and bandwidth as in the real network. This ensures that the state and action spaces remain statistically identical. Then, the wireless channel model is used to compute the RSRP between the BSs and users, and the user mobility model generates unique trajectories for each user in the environment. As users keep entering and leaving this area, the total number of users, $|\mathcal{U}|$, evolves over a long-time scale.
@@ V-B Parallel DT Framework
Specifically, the evolution of $|\mathcal{U}|$ in an environment occurs over a long-time scale (hundreds of time slots), while sample collection and batch training of $\pi_{\boldsymbol{\theta}}$ are performed over just one or a few time slots. As a result, the number of users in each training batch remains relatively stable or fluctuates slightly. This strong correlation of training samples on $|\mathcal{U}|$ would cause non-stationarity in DRL training, resulting in a suboptimal policy that focuses primarily on the current state, as schematized in Fig. 5.
@@
One potential solution to this issue in a single environment is to extend both the training interval and batch size, ensuring the batch training occurs only after $|\mathcal{U}|$ has sufficiently evolved. However, this approach would significantly reduce training efficiency. In multi-agent DRL systems, $\pi_{\boldsymbol{\theta}}$ can only undergo a limited number of gradient descents before being redeployed for further exploration. This is because the state and transition probabilities of the MDP for each agent are significantly affected by the actions of other agents. In the environment studied in this paper, this effect is observed in the strong relationship between actions of other agents under policy $\pi_{\boldsymbol{\theta}}$ and the distribution and transition probabilities of $\boldsymbol{l}[n]$ within the state $\mathbf{s}_{i}[n]$. While extending the training interval can mitigate this correlation, only a small portion of the collected samples would be effectively utilized for training, thus reducing the training efficiency. For instance, in our experiments, the traversal of the number of users from 100 to 400 requires 20,000 environment steps, generating $5\times 10^{6}$ training samples, which are sufficient for DRL agents to converge. However, with an extended training interval, these samples would only support a few gradient descents, resulting in most of the samples being wasted. This low training efficiency is impractical for DRL training in both real and DT environments.
@@
In summary, the proposed parallel DT-driven DRL method consists of the distributed DRL method, the LVM-enhanced DT, and the parallel DT framework. It addresses high trial-and-error costs and poor physical network performance by training in virtual DT environments. It also alleviates the strong correlation and non-stationarity in the real environment through parallelization, which therefore improves training efficiency and the generalization ability of DRL models. Note that the LVM-enhanced DT and parallel DT framework are versatile and applicable for other DT-driven network optimization tasks, particularly mobility-sensitive ones, such as wireless resource management [46] and task offloading in mobile edge computing [47].
@@ VI Performance Evaluation
In this section, we assess the efficacy of our proposed parallel DT-driven DRL method. First, we examine the established DT’s fidelity to the real environment and its reliability in the DRL training. Then, a comprehensive evaluation of the proposed parallel DT-DRL method is conducted, along with a comparison against the DRL method trained directly in the real environment.
@@ VI-B Dataset and Experiment setting
Map2Traj is trained using a real-world vehicle trajectory dataset from Xi’an, China, recorded in 2016 [48], alongside the OpenStreetMap dataset of the same year. These datasets are sourced from the ChinaGEOSS Data Sharing Network. The trajectories are all within the latitude range of 34.21 to 34.28 and the longitude range of 108.912 to 108.996. To avoid data leakage in the zero-shot trajectory generation, the training set is limited to longitudes between 108.912 and 108.974, while the test set extends from 108.974 to 108.996. Notice that the correlation between the street map and trajectory, as shown in Fig. 2, remains consistent under transformations such as rotation and reflection. This inherent property can be leveraged for data augmentation. We randomly rotate and flip both street maps and trajectory data in training to enhance the generalization capability of the Map2Traj model.
@@ VI-C Evaluation Metrics
The metrics for spatial distribution consist of cosine similarity and Wasserstein distance [53]. Cosine similarity is a widely used measure of similarity between two vectors. While it reflects the similarity between probability distributions, it falls short in expressing the spatial correlation between adjacent blocks in two-dimensional (2D) distributions. To address this limitation, we introduce the Wasserstein distance, which quantifies the minimal cost required to transform one probability distribution into another through an optimal transport plan. In our experiment, the Wasserstein distance measures the effort required to transform the spatial distribution of generated trajectories into that of real trajectories. However, computing Wasserstein distance for 2D distributions entails solving a complex high-dimensional linear programming problem, which is computationally impractical for 192$\times$192 arrays. Consequently, we use the sliced Wasserstein distance [54] as an alternative.
@@ VI-D Numerical Results
We select the area mentioned in V-A and Fig. 2 as the test area. Fig. 8 displays the generated trajectories by all methods alongside corresponding heatmaps. Traditional random mobility models, i.e., RWP and GM, result in chaotic trajectories and heatmaps that bear no resemblance to real-world patterns. While map-restricted models, M-RWP and M-GM, show some similarity in trajectory patterns to real ones, they still fall short in spatial distribution similarity due to the absence of a learning mechanism. As expected, DiffTraj demonstrates high similarity to real trajectories and heatmaps because it is trained directly on real data. In contrast, our proposed Map2Traj, even in a zero-shot scenario, produces results comparable to real trajectories and even surpasses DiffTraj in trajectory details.
(a) RWP
(b) GM
(c) M-RWP
(d) M-GM
(e) DiffTraj
(f) Map2Traj
@@
It is encouraging to see that our zero-shot Map2Traj model outperforms the area-specific DiffTraj. The advantage of Map2Traj lies in the continuous guidance from street maps throughout the denoising process, while DiffTraj only involves some trip information as a condition. On the other hand, DiffTraj fixes the trajectories into uniform shapes through sampling, instead of transforming them into images, potentially leading to information loss. However, it is crucial to acknowledge that existing learning-based methods, including DiffTraj, have the potential to outmatch Map2Traj given a sufficiently large training dataset, a wider and deeper network structure, or in some specific test scenarios. The primary contribution of Map2Traj is the zero-shot trajectory generation ability, rather than merely surpassing existing area-specific trajectory generation techniques.
Figure 9: DRL performance in the training environment.
@@
Figure 10: DRL performance after deployment in a real environment.
@@
As illustrated in Fig. 11, both similarities in spatial distribution and trajectory pattern are positively correlated with the performance of DRL agents. This finding verifies our analysis of the prerequisite for constructing faithful DT for network optimization, as discussed in Section V-A.
@@
Figure 13: DRL performance in the training environment
@@
Figure 14: DRL performance in the real environment
@@ VII Conclusion
In this paper, we have proposed a parallel DT-driven DRL method designed for user association and load balancing in dynamic wireless networks. To accommodate varying user numbers, we have introduced a distributed DRL strategy with an enhanced PPO network structure to speed up convergence. To address high trial-and-error costs in real-world training and poor physical network performance before convergence, we have developed an LVM-enhanced DT for wireless networks, featuring a zero-shot generative user mobility model named Map2Traj. DRL agents undergo training within this DT environment, thereby avoiding direct interaction with the physical network. Furthermore, a parallel DT framework has been developed to alleviate the strong correlation and non-stationarity in the real environment through parallelization, which therefore improves training efficiency and the generalization ability of DRL models. The proposed parallel DT-driven DRL method is versatile and applicable for other DT-driven network optimization tasks, such as wireless resource management and task offloading in mobile edge computing.